OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies inversely as x and y=7 when x=12. The variation constant is ??

OpenStudy (anonymous):

...this is SO beyond me!

OpenStudy (saifoo.khan):

y = k/x

OpenStudy (anonymous):

y = k/x, 7 = k/12, k=84

OpenStudy (saifoo.khan):

7 = k/12 k = 7*12 k = 84

OpenStudy (anonymous):

First I need the variation constant??

OpenStudy (saifoo.khan):

Always!

OpenStudy (anonymous):

I have no idea what that is :(

OpenStudy (anonymous):

y+k/12??

OpenStudy (anonymous):

=

OpenStudy (saifoo.khan):

umm.

OpenStudy (saifoo.khan):

\[y = \frac{k}{x}\]

OpenStudy (saifoo.khan):

like that^